OpenStudy (anonymous):

If -nx2 + tx + c = 0, what is x equal to?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

You could use the quadratic formula, to solve for x in terms of n, t and c.

OpenStudy (anonymous):

idk how

OpenStudy (agent0smith):

if you have \[\large ax^2 + bx + c = 0\] then \[\large x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] All you need to do is identify what a, b and c are here.

OpenStudy (anonymous):